find the volume of a cone in ft3 with the following dimensions:

Diameter =0.00235km

height = 12345 mm

1 m= 3285ft


1
Expert's answer
2021-10-13T09:07:14-0400

We have to use all the conversions and the information given to find the volume expressed in ft3:


"V=\frac{1}{3}\pi r^2h=\frac{1}{3}\pi (\frac{d}{2})^2h
\\ V=\frac{1}{3}\pi (\frac{0.00235\,km}{2} \times \frac{1000\,m}{1\,km} \times \frac{3285\,ft}{1\,m})^2(12345\,mm\times \frac{1\,m}{1000\,mm} \times \frac{3285\,ft}{1\,m})
\\ V=\frac{1}{3}\pi (3859.875\,ft)^2(40553.325\,ft)
\\ V= 6.32705\times 10^{11}\,{ft}^3"


In conclusion, we were able to find that Vcone = 6.327 X 1011 ft3.

Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.